# CMRNext: Camera to LiDAR Matching in the Wild for Localization and Extrinsic Calibration

### v3

@@ III-E Iterative Refinement
Due to the incorrect initial position , the LiDAR-Image and the camera image can have a small overlapping Field of View (FoV), thus sharing only a few points that can be matched by our method. We therefore employ an iterative refinement approach, in which multiple instances of CMRNext are independently trained using different initial error ranges. During inference, the pair (, ) is fed to the instance of CMRNext trained with the highest initial error range and the predicted pose is used to generate a new synthesized LiDAR-image . The latter is subsequently fed to the second instance of CMRNext, which is trained with a lower initial error range. This process can be repeated multiple times, using instances of CMRNext trained with increasingly small error ranges, iteratively reducing the estimated localization error.
@@ IV Experimental Results
In this section, we will detail and analyze the extensive experimental evaluations we performed to validate our novel CMRNext. First, we present the datasets we used to train and evaluate our approach, followed by the training and implementation details. We then report qualitative and quantitative results on two different tasks: monocular localization in LiDAR maps and extrinsic camera-LiDAR calibration. Although the two tasks are similar in nature, they tackle two different use cases and they present different and complementary challenges. Finally, we present multiple ablation studies aimed at validating the architectural choices we made when designing CMRNext. Unless otherwise specified, to mitigate the effect of the random initial error added to the ground truth poses, all experiments were performed using three different random seeds and the results were averaged.
@@ IV-A1 KITTI
The KITTI dataset [13] was recorded in Karlsruhe, Germany, and includes both urban and interurban scenarios. Accurate localization and semantic information for the LiDAR point clouds are additionally provided for the “odometry” sequences in the SemanticKITTI dataset [54]. The recording vehicle was equipped with two forward-facing RGB cameras in a stereo configuration that provides images ranging from to pixels at 10 FPS. It is also equipped with a Velodyne HDL-64E that we used to generate the LiDAR-maps. We include the left camera images from the odometry sequences 03, 05, 06, 07, 08, and 09 in the training set (total of frames) and we use the sequence 00 for evaluation ( frames).
@@ IV-A4 In-House Datasets
We additionally recorded three in-house datasets by driving around the city of Freiburg, Germany, and around our university campus using three different robotic platforms, a Self-driving Perception Car, a Quadruped robot, and a quadcopter UAV. The three platforms are depicted in the lower box of Figure 1. All platforms are equipped with an ouster OS1-128 LiDAR and a forward-facing RGB camera. Resolution and field of view of the cameras mounted on the three platforms are different and are reported in Table I. The in-house datasets are used to evaluate the generalization ability of our approach to different cities and sensor setups and therefore are not included in the training set. In particular, the Self-driving Perception Car dataset is used to evaluate the generalization ability on the monocular localization task, while all three datasets are used to evaluate the generalization ability on the extrinsic calibration task, as the sensors on the Quadruped and UAV platforms are not time-synchronized.
@@ IV-B Training Details
Unless otherwise specified, we train CMRNext on four NVIDIA GeForce RTX 3090 with a total batch size of 16, using the Adam optimizer with a base learning rate of and a weight decay of . We use a OneCycle learning rate scheduler and we train the network in two stages. In the first stage, we train the network without uncertainty estimation for 150 epochs using the original MAE-based loss function from RAFT [48]. In the second stage, we fine-tune the network with uncertainty estimation for 50 epochs using the loss function defined in Equation 15. We use frequencies for the Fourier feature mapping module and we set in Equation 15.
@@ IV-D Training on Multiple Datasets
We train CMRNext by combining training samples from KITTI, Argoverse, and Pandaset datasets. Training a CNN on multiple diverse datasets creates certain challenges. First, the different cardinality of the three training datasets (, , and , respectively) might lead the network to perform better on one dataset than the other. To overcome this problem, we randomly sampled a subset of KITTI and Argoverse at the beginning of every epoch to have the same number of samples as Pandaset. As the subset is sampled every epoch, the network will eventually process every sample from the datasets.
Moreover, the three datasets have cameras with very different fields of view and resolution. To address this issue, we resize the images so to have the same resolution. One straightforward way to accomplish this would be to just crop the Argoverse and Pandaset images. However, this would yield images with a very narrow field of view, making the matching between the RGB and LiDAR-image increasingly difficult. Therefore, we first downsample the Argoverse and Pandaset images to half the original resolution, and then randomly crop images from all datasets to pixels. We perform this random cropping at runtime during training in order to have different crop positions for the same image at different epochs. We generate both the LiDAR-image and the ground truth displacements at the original resolution, and then downsample and crop them accordingly. Moreover, we also halve the pixel displacements during the downsampling operation.
@@ IV-E Initial Pose Sampling and Data Augmentation
To improve the generalization ability of our approach, we employ a data augmentation scheme. First, we apply color jittering to the images by randomly changing the contrast, saturation, and brightness. Subsequently, we randomly horizontally mirror the imagesand we modify the camera calibration accordingly. We then randomly rotate the images in the range [, ]. Finally, we transform the LiDAR point cloud to reflect these changes before generating the LiDAR-image.
@@
Randomly draw a subset of the KITTI and Argoverse datasets and shuffle them with samples from Pandaset
@@ IV-F Inference
During inference, we process one image at a time and since the network architecture is fully convolutional, we do not resize the images of different datasets to have the same resolution. Therefore, we feed the whole image to CMRNext without any cropping. However, we still downsample the images of the Argoverse, Pandaset, and our in-house datasets to have a similar field of view as the images used for training. Once we have the set of 2D-3D correspondences predicted by the network, we apply EPnP+RANSAC to estimate the pose of the camera with respect to the map, as detailed in Section III-B. We repeat this whole process three times using three specialized instances of CMRNext to iteratively refine the estimation.
@@ IV-G Monocular Localization
| KITTI | Argoverse | Pandaset | Freiburg - Car | |||||
| Transl. [cm] | Rot. [°] | Transl. [cm] | Rot. [°] | Transl. [cm] | Rot. [°] | Transl. [cm] | Rot. [°] | |
| Initial Pose | 188 | 9.90 | 187 | 9.87 | 193 | 9.80 | 188 | 9.87 |
| I2P-Net [61] | 7 | 0.67 | - | - | - | - | - | - |
| I2D-Loc [62] | 8 | 0.30 | - | - | - | - | - | - |
| CMRNext (ours) | 6.21 | 0.23 | 7.51 | 0.16 | 7.18 | 0.11 | 10.11 | 0.41 |
We evaluate the monocular localization performance of CMRNext on the KITTI, Argoverse, Pandaset, and our in-house datasets. We compare our approach with the state-of-the-art camera-dependent methods CMRNet [20], HyperMap [21], and I2P-Net [61], as well as camera-agnostic methods CMRNet++ [26] and I2D-Loc [62].
Camera-dependent methods, such as CMRNet, I2P-Net, and HyperMap, can only be evaluated on the same dataset they were trained on and cannot generalize to different sensor setups. Therefore, for these methods, a different model is trained for each dataset. HyperMap and I2P-Net do not provide a publicly available implementation and therefore we report the results from their respective papers [21, 61]. For CMRNet, we used the publicly available implementation111https://github.com/cattaneod/CMRNet and we trained a model for each dataset. As our in-house Freiburg dataset is not included in the training set, but only used for evaluating the generalization ability, no camera-dependent method can be evaluated on this dataset. On the other hand, camera-agnostic methods, such as I2D-Loc, CMRNet++, and our proposed CMRNext, can be trained on multiple datasets simultaneously and the same model can be used to evaluate the performance on all four datasets. Although the code I2D-Loc is publicly available222https://github.com/EasonChen99/I2D-Loc, only the training on the KITTI dataset is provided. While trying to train I2D-Loc on multiple datasets, we encountered several issues and we were unable to reproduce the results reported in the original paper. Therefore, we report the results from the I2D-Loc paper [62] on KITTI and Argoverse and we evaluated the pretrained model provided by the authors on Pandaset and Freiburg. All methods except for I2P-Net use an iterative refinement approach, where multiple networks are trained with different initial error ranges. To provide a fair comparison, for these methods, we only report the results of the first iteration, as the number and error ranges of successive iterations vary between methods.
@@
Notably, CMRNext is able to effectively generalize to different sensor setups and different cities than those used for training, while existing camera-agnostic methods struggle to do so. As previously mentioned, camera-dependent methods are tied to a specific sensor setup and therefore they cannot be evaluated on unseen datasets. In particular, our method achieves a median translation error of and a median rotation error of on our in-house dataset, compared to and achieved by the second-best method.
@@ IV-G1 Iterative Refinement
We additionally compare the performance of CMRNext with the best performing camera-dependent (I2P-Net) and camera-agnostic (I2D-Loc) methods. For this comparison, the final iteration of all methods is reported. Both I2P-Net and I2D-Loc report these results only on the KITTI dataset. As shown in Table III, CMRNext outperforms all methods, setting the new state-of-the-art for monocular localization in LiDAR maps. Specifically, CMRNext achieves a median translation error of , compared to the second-best method I2P-Net, which achieves a median translation error of . Moreover, CMRNext achieves a median rotation error of , compared to the second best method I2D-Loc, which achieves a median rotation error of .
In Figure 5 we show qualitative results of CMRNext for the monocular localization task on the KITTI, Argoverse, Pandaset, and Freiburg-Car datasets. The first column shows the LiDAR-map projection in the initial pose overlaid with the camera image, the second column shows the projection in the ground truth camera pose , and the third column shows the projection in the pose predicted by CMRNext. We observe that our method is able to accurately localize the camera in the LiDAR-map, even in challenging scenarios with dynamic objects that are not present in the map. The resulting LiDAR-map projections align well with the camera images.
@@ IV-H Ablation Study
In Table IV we evaluate the performance of CMRNext using different optical flow networks. In particular, we evaluate established optical flow networks, namely PWCNet [39], MaskFlowNet [63] and RAFT [48], as well as the recent state-of-the-art methods GMFlow [49] and FlowFormer [64]. Unsurprisingly, RAFT outperforms previous methods PWCNet and MaskFlowNet by a large margin. However, we noticed that both transformer-based methods GMFlow and FlowFormer underperformed compared to RAFT, although they achieved better results on optical flow benchmarks. We hypothesize that this is due to the fact that networks based on Vision Transformers (ViT) lack intrinsic inductive bias [65] and thus require a large amount of data to train. The vast majority of methods for optical flow estimation are pretrained on large-scale synthetic datasets, such as FlyingChairs and FlyingThings, and therefore ViT-based methods are able to learn this inductive bias. For our cross-modal task, however, such large-scale datasets are not available and therefore the CNN-based method RAFT achieves the best performance among all the evaluated networks. Moreover, the positional encoding used in transformers is sensitive to the shape of the input, and during inference images with the same resolution as the ones used during training should be used for optimal performance. The authors of FlowFormer [64] suggested a tiling strategy, where the input image is split into multiple tiles, each tile is then independently processed by the network and the resulting optical flow fields are combined together. While this strategy substantially improves the performance of FlowFormer (see Table IV), it requires multiple forward passes of the network, drastically increasing the computational cost.
@@
Next, in the second set of rows, we verified that the occlusion filter considerably reduces the localization error of CMRNext, especially on the Argoverse and Pandaset datasets. The camera images of these datasets have a larger field of view and higher resolution than those of the KITTI dataset and thus occluded map points are more likely to be projected in the LiDAR images if no occlusion filter is applied, negatively impacting the performance.
@@ IV-I LiDAR-Camera Calibration
We extended our method to the task of LiDAR-camera extrinsic calibration. Although this task shares some similarities with monocular localization in LiDAR maps, there are some key differences. First, in the calibration task, the camera image is matched against a single LiDAR scan recorded at the same time, therefore occlusions and dynamic objects do not hinder the matching process, as both sensors observe the same scene. For this reason, we did not use the occlusion filter for all the experiments reported in this section. Second, for a specific LiDAR-camera pair, the projection of the LiDAR points in the camera frame using the ground truth extrinsic calibration would contain some recognizable pattern induced by the LiDAR sensor specification and placement. For example, the horizon line of the LiDAR scan would be projected on the same row of pixels in the camera image and the typical ring pattern of mechanical LiDAR sensors would appear very similar on all images, as illustrated in Figure 6. Networks that are trained on a single LiDAR-camera pair could learn to exploit these patterns and thus overfit to the specific extrinsic calibration of the training dataset. This is especially true for camera-dependent methods, as the only way to include multiple extrinsic calibrations without changing the camera is to physically move the sensors, drastically increasing the effort required to collect a training dataset. Most of the existing DNN-based methods for LiDAR-camera calibration [38, 61, 35, 36, 59] are trained and evaluated on the same extrinsic calibration. We believe that this evaluation protocol is misleading, as a network that always predicts the same extrinsic parameters cannot be used in any practical application. Therefore, we evaluate all methods on LiDAR-camera pairs that were not included in the training set. In particular, we used the right camera of the KITTI odometry dataset for evaluation, while only the left camera was included in the training.
@@
We compared the performance of CMRNext on the LiDAR-camera extrinsic calibration task with our previous methods CMRNet [20], as well as with the state-of-the-art methods RGGNet [37] and LCCNet [38]. For the latter methods, we used the pretrained models provided by the respective authors333https://github.com/LvXudong-HIT/LCCNet444https://github.com/KleinYuan/RGGNet. Other existing methods [61, 36, 41, 59] do not provide a publicly available implementation and therefore we were unable to evaluate them on the right camera of the KITTI dataset, as discussed above. In Table VI, we report the results of CMRNet, LCCNet, and CMRNext after the iterative refinement (five iterations for CMRNet and LCCNet, and three iterations for CMRNext). The results show that although LCCNet achieves the lowest translation error when evaluated on the same camera used for training, it fails to generalize to a different camera, still predicting the extrinsic for the left camera when evaluated on the right one. We observe a similar behavior for our previous work CMRNet, as both CMRNet and LCCNet are camera-dependent and thus more prone to overfitting to the specific extrinsic. Our proposed CMRNext, on the other hand, outperforms existing methods on the right camera, with a median translation error of and a median rotation error of , while at the same time achieving the lowest rotation error and comparable translation error on the left camera.
@@
RGGNet [37] was trained on raw sequences of the KITTI dataset, instead of the odometry sequences and with an initial error up to [ , ]. To provide a fair comparison with RGGNet, we trained a single instance of CMRNext without iterative refinement using the same training set used by RGGNet and we evaluated on both left and right cameras of the T1 and T3 testing datasets defined in [37]. The testing datasets T2a and T2b contain sequences that were included in the training set and therefore do not provide a fair evaluation, hence we excluded them from the evaluation. -RegNet is the reimplementation of RegNet [35] by the authors of RGGNet. In addition to the median rotation and translation error, following [37] for this experiment we also report the mean error (MSEE) and mean re-calibration rate (MRR), defined as:

### v4

@@ III-E Iterative Refinement
Due to the incorrect initial position , the LiDAR-Image and the camera image can have a small overlapping Field of View (FoV), thus sharing only a few points that can be matched by our method. We, therefore, employ an iterative refinement approach in which multiple instances of CMRNext are independently trained using different initial error ranges. During inference, the pair (, ) is fed to the instance of CMRNext trained with the highest initial error range, and the predicted pose is used to generate a new synthesized LiDAR-image . The latter is subsequently fed to the second instance of CMRNext, which is trained with a lower initial error range. This process can be repeated multiple times, using instances of CMRNext trained with increasingly small error ranges, iteratively reducing the estimated localization error.
@@ IV Experimental Results
In this section, we will detail and analyze the extensive experimental evaluations we performed to validate our novel CMRNext. First, we present the datasets we used to train and evaluate our approach, followed by the training and implementation details. We then report qualitative and quantitative results on two different tasks: monocular localization in LiDAR maps and extrinsic camera-LiDAR calibration. Although the two tasks are similar in nature, they tackle two different use cases, and they present different and complementary challenges. Finally, we present multiple ablation studies aimed at validating the architectural choices we made when designing CMRNext. Unless otherwise specified, to mitigate the effect of the random initial error added to the ground truth poses, all experiments were performed using three different random seeds, and the results were averaged.
@@ IV-A1 KITTI
The KITTI dataset [13] was recorded in Karlsruhe, Germany, and includes both urban and interurban scenarios. Accurate localization and semantic information for the LiDAR point clouds are additionally provided for the “odometry” sequences in the SemanticKITTI dataset [57]. The recording vehicle was equipped with two forward-facing RGB cameras in a stereo configuration that provides images ranging from to pixels at 10 FPS. It is also equipped with a Velodyne HDL-64E that we used to generate the LiDAR-maps. We include the left camera images from the odometry sequences 03, 05, 06, 07, 08, and 09 in the training set (total of frames), and we use the sequence 00 for evaluation ( frames).
@@ IV-A4 In-House Datasets
[…]
@@ IV-B Training Details
Unless otherwise specified, we train CMRNext on four NVIDIA GeForce RTX 3090 with a total batch size of 16, using the Adam optimizer with a base learning rate of and a weight decay of . We use a OneCycle learning rate scheduler and we train the network in two stages. In the first stage, we train the network without uncertainty estimation for 150 epochs using the original MAE-based loss function from RAFT [51]. In the second stage, we fine-tune the network with uncertainty estimation for 50 epochs using the loss function defined in Equation 15. We use frequencies for the Fourier feature mapping module in Equation 10 and we set in Equation 15.
@@ IV-D Training on Multiple Datasets
We train CMRNext by combining training samples from KITTI, Argoverse, and Pandaset datasets. Training a CNN on multiple diverse datasets creates certain challenges. First, the different cardinality of the three training datasets (, , and , respectively) might lead the network to perform better on one dataset than the other. To overcome this problem, we randomly sampled a subset of Argoverse and Pandaset at the beginning of every epoch to have the same number of samples as KITTI. As the subset is sampled every epoch, the network will eventually process every sample from the datasets.
Moreover, the three datasets have cameras with very different fields of view and resolution. To address this issue, we resize the images so to have the same resolution. One straightforward way to accomplish this would be to just crop the Argoverse and Pandaset images. However, this would yield images with a very narrow field of view, making the matching between the RGB and LiDAR-image increasingly difficult. Therefore, we first downsample the Argoverse and Pandaset images to half the original resolution and then randomly crop images from all datasets to pixels. We perform this random cropping at runtime during training in order to have different crop positions for the same image at different epochs. Since we generate both the LiDAR-image and the ground truth displacements at the original resolution, and only after we downsample and crop them, we don’t need to update the intrinsic parameters of the camera. Moreover, we also halve the pixel displacements during the downsampling operation.
@@ IV-E Initial Pose Sampling and Data Augmentation
To improve the generalization ability of our approach, we employ a data augmentation scheme. First, we apply color jittering to the images by randomly changing the contrast, saturation, and brightness. Subsequently, we randomly horizontally mirror the images and we modify the camera calibration by updating the principal point as follows: , where is the width of the image. We then randomly rotate the images in the range [, ]. Finally, we transform the LiDAR point cloud to reflect these changes before generating the LiDAR-image. The LiDAR point clouds are expressed with the camera pose as the origin, and with reference frame X-forward, Y-left, and Z-up. To reflect the horizontal mirror of the image, we multiply the Y-coordinates by , and we rotate around the X-axis by the same angle used to rotate the image.
@@
Randomly draw a subset of the Argoverse and Pandaset datasets and shuffle them with samples from KITTI
@@ IV-F Inference
During inference, we process one image at a time, and since the network architecture is fully convolutional, we do not crop the images of different datasets to have the same resolution. Therefore, we feed the whole image to CMRNext. However, we still downsample the images of the Argoverse, Pandaset, and our in-house datasets to have a similar field of view as the images used for training, and we successively upsample the displacement flow predicted by CMRNext to the original resolution and scale it accordingly. Once we have the set of 2D-3D correspondences predicted by the network, we apply EPnP+RANSAC to estimate the pose of the camera with respect to the map, as detailed in Section III-B. We repeat this whole process three times using three specialized instances of CMRNext to iteratively refine the estimation.
@@ IV-G Monocular Localization
| KITTI | Argoverse | Pandaset | Freiburg - Car | ||||||
| Initial Pose | 188 | 9.90 | 187 | 9.87 | 193 | 9.80 | 188 | 9.87 | |
Single
Iteration |
HyperMap [22] | 48 | 1.42 | - | - | - | - | - | - |
| BEVLoc [63] | 39 | 1.28 | - | - | - | - | - | - | |
| I2P-Net [64] | 7 | 0.67 | - | - | - | - | - | - | |
Iterative
Refinement |
CMRNet [21] | 27 | 1.07 | - | - | - | - | - | - |
| PSNet [62] | 25 | 0.63 | - | - | - | - | - | - | |
| POET [23] | 20 | 0.79 | - | - | - | - | - | - | |
| I2D-Loc [65] | 8 | 0.30 | - | - | - | - | - | - | |
| CMRNet++ [27] | 14 | 0.43 | 25 | 0.45 | - | - | - | - | |
| CMRNext (same range) | 6.33 | 0.23 | 19.09 | 0.31 | 8.39 | 0.13 | 17.73 | 0.27 | |
| CMRNext (ours) | 6.21 | 0.23 | 7.51 | 0.16 | 7.18 | 0.11 | 12.94 | 0.28 | |
We evaluate the monocular localization performance of CMRNext on the KITTI, Argoverse, Pandaset, and our in-house datasets. We compare our approach with the state-of-the-art camera-dependent methods CMRNet [21], HyperMap [22], PSNet [62], BEVLOC [63], POET [23], and I2P-Net [64], as well as camera-agnostic methods CMRNet++ [27] and I2D-Loc [65].
Camera-dependent methods can only be evaluated on the same dataset they were trained on and cannot generalize to different sensor setups. Therefore, for these methods, a different model is trained for each dataset. HyperMap, PSNet, BEVLOC, POET, and I2P-Net do not provide a publicly available implementation, and therefore, we report the results from their respective papers [22, 62, 63, 23, 64]. For CMRNet, we used the publicly available implementation222https://github.com/cattaneod/CMRNet and we trained a model for each dataset. As our in-house Freiburg dataset is not included in the training set, but only used for evaluating the generalization ability, no camera-dependent method can be evaluated on this dataset. On the other hand, camera-agnostic methods, such as I2D-Loc, CMRNet++, and our proposed CMRNext, can be trained on multiple datasets simultaneously, and the same model can be used to evaluate the performance on all four datasets. Although the code I2D-Loc is publicly available333https://github.com/EasonChen99/I2D-Loc, only the training on the KITTI dataset is provided. While trying to train I2D-Loc on multiple datasets, we encountered several issues, and we were unable to reproduce the results reported in the original paper. Therefore, we report the results from the I2D-Loc paper [65] on KITTI and Argoverse, and we evaluated the pretrained model provided by the authors on Pandaset and Freiburg. All methods except for I2P-Net use an iterative refinement approach, where multiple networks are trained with different initial error ranges. To provide a fair comparison, for these methods, we only report the results of the first iteration, as the number and error ranges of successive iterations vary between methods.
@@
Notably, CMRNext is able to effectively generalize to different sensor setups and different cities than those used for training, while existing camera-agnostic methods struggle to do so. As previously mentioned, camera-dependent methods are tied to a specific sensor setup, and therefore, they cannot be evaluated on unseen datasets. In particular, our method achieves a median translation error of and a median rotation error of on our in-house dataset, compared to and achieved by the second-best method.
@@ IV-G1 Iterative Refinement
We additionally compare the performance of CMRNext with camera-dependent and camera-agnostic methods. While HyperMap, BEVLoc, and I2P-Net do not provide iterative refinement, all the other methods use three instances trained with the following error ranges: [, ], [, ], and [, ]. For a fair comparison, we additionally evaluate our method using the same error ranges used by the baselines, denoted as CMRNext (same range). For this experiment, we only report the results after the final iteration of all methods. Most baselines report these results only on the KITTI dataset. As shown in Table III, CMRNext (same range) outperforms all methods, setting the new state-of-the-art for monocular localization in LiDAR maps. By using our optimized initial error ranges in CMRNext (ours), the translation and rotation errors further decrease on most datasets. Specifically, CMRNext achieves a median translation error of , compared to the second-best method I2P-Net, which achieves a median translation error of . Moreover, CMRNext achieves a median rotation error of , compared to the second best method, I2D-Loc which achieves a median rotation error of .
In Figure 5, we present qualitative results of CMRNext for the monocular localization task on the KITTI, Argoverse, Pandaset, and Freiburg-Car datasets. The first column shows the LiDAR-map projection in the initial pose overlaid with the camera image, the second column shows the projection in the ground truth camera pose , and the third column shows the projection in the pose predicted by CMRNext. We observe that our method is able to accurately localize the camera in the LiDAR-map, even in challenging scenarios with dynamic objects that are not present in the map. The resulting LiDAR-map projections align well with the camera images. Finally, in Figure 6, we report the rotation and translation errors of CMRNext on the sequence 00 of the KITTI dataset, alongside images from three scenes where CMRNext achieved higher errors. We observed that the highest errors typically occur when the car is making a turn or when the environment is predominantly filled with trees and lacks structural elements, making feature matching particularly challenging.
@@ IV-H Ablation Study
In Table IV we evaluate the performance of CMRNext using different optical flow networks. In particular, we evaluate established optical flow networks, namely PWCNet [40], MaskFlowNet [66] and RAFT [51], as well as the recent state-of-the-art methods GMFlow [52] and FlowFormer [67]. Unsurprisingly, RAFT outperforms previous methods PWCNet and MaskFlowNet by a large margin. However, we noticed that both transformer-based methods GMFlow and FlowFormer underperformed compared to RAFT, although they achieved better results on optical flow benchmarks. We hypothesize that this is due to the fact that networks based on Vision Transformers (ViT) lack intrinsic inductive bias [68] and thus require a large amount of data to train. The vast majority of methods for optical flow estimation are pretrained on large-scale synthetic datasets, such as FlyingChairs and FlyingThings, and therefore ViT-based methods are able to learn this inductive bias. For our cross-modal task, however, such large-scale datasets are not available, and therefore, the CNN-based method RAFT achieves the best performance among all the evaluated networks. Moreover, the positional encoding used in transformers is sensitive to the shape of the input, and during inference, images with the same resolution as the ones used during training should be used for optimal performance. The authors of FlowFormer [67] suggested a tiling strategy where the input image is split into multiple tiles, each tile is then independently processed by the network, and the resulting optical flow fields are combined together. While this strategy substantially improves the performance of FlowFormer (see Table IV), it requires multiple forward passes of the network, drastically increasing the computational cost.
@@
Next, in the second set of rows, we verified that the occlusion filter considerably reduces the localization error of CMRNext, especially on the Argoverse and Pandaset datasets. The camera images of these datasets have a larger field of view and higher resolution than those of the KITTI dataset, and thus, occluded map points are more likely to be projected in the LiDAR images if no occlusion filter is applied, negatively impacting the performance.
@@ IV-I LiDAR-Camera Calibration
We extended our method to the task of LiDAR-camera extrinsic calibration. Although this task shares some similarities with monocular localization in LiDAR maps, there are some key differences. First, in the calibration task, the camera image is matched against a single LiDAR scan recorded at the same time, therefore occlusions and dynamic objects do not hinder the matching process, as both sensors observe the same scene. For this reason, we did not use the occlusion filter for all the experiments reported in this section. Second, for a specific LiDAR-camera pair, the projection of the LiDAR points in the camera frame using the ground truth extrinsic calibration would contain some recognizable pattern induced by the LiDAR sensor specification and placement. For example, the horizon line of the LiDAR scan would be projected on the same row of pixels in the camera image, and the typical ring pattern of mechanical LiDAR sensors would appear very similar on all images, as illustrated in Figure 7. Networks that are trained on a single LiDAR-camera pair could learn to exploit these patterns and thus overfit to the specific extrinsic calibration of the training dataset. This is especially true for camera-dependent methods, as the only way to include multiple extrinsic calibrations without changing the camera is to physically move the sensors, drastically increasing the effort required to collect a training dataset. Most of the existing DNN-based methods for LiDAR-camera calibration [39, 64, 36, 37, 62] are trained and evaluated on the same extrinsic calibration. We believe that this evaluation protocol is misleading, as a network that always predicts the same extrinsic parameters cannot be used in any practical application. Therefore, we evaluate all methods on LiDAR-camera pairs that were not included in the training set. In particular, we used the right camera of the KITTI odometry dataset for evaluation, while only the left camera was included in the training.
@@
We compared the performance of CMRNext on the LiDAR-camera extrinsic calibration task with our previous methods CMRNet [21], as well as with the state-of-the-art methods RGGNet [38] and LCCNet [39]. For the latter methods, we used the pretrained models provided by the respective authors444https://github.com/LvXudong-HIT/LCCNet555https://github.com/KleinYuan/RGGNet. Other existing methods [64, 37, 44, 62] do not provide a publicly available implementation and therefore we were unable to evaluate them on the right camera of the KITTI dataset, as discussed above. In Table VI, we report the results of CMRNet, LCCNet, and CMRNext after the iterative refinement (five iterations for CMRNet and LCCNet, and three iterations for CMRNext). The results show that although LCCNet achieves the lowest translation error when evaluated on the same camera used for training, it fails to generalize to a different camera, still predicting the extrinsic for the left camera when evaluated on the right one. We observe a similar behavior for our previous work CMRNet, as both CMRNet and LCCNet are camera-dependent and thus more prone to overfitting to the specific extrinsic. Our proposed CMRNext, on the other hand, outperforms existing methods on the right camera, with a median translation error of and a median rotation error of , while at the same time achieving the lowest rotation error and comparable translation error on the left camera.
@@
RGGNet [38] was trained on raw sequences of the KITTI dataset, instead of the odometry sequences and with an initial error up to [ , ]. To provide a fair comparison with RGGNet, we trained a single instance of CMRNext without iterative refinement using the same training set used by RGGNet and we evaluated on both left and right cameras of the T1 and T3 testing datasets defined in [38]. The testing datasets T2a and T2b contain sequences that were included in the training set and therefore do not provide a fair evaluation, hence we excluded them from the evaluation. -RegNet is the reimplementation of RegNet [36] by the authors of RGGNet. In addition to the median rotation and translation error, following [38] for this experiment, we also report the mean error (MSEE) and mean re-calibration rate (MRR), defined as: